OpenStudy (anonymous):

(5) what is value of the expression when a=5, b=4 and c=6? 4a/2b-c A. 10 B. 5 C. 9 D. 2

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

4(5) / 2(4) - 6 simplify it pls

OpenStudy (anonymous):

Hello :)

OpenStudy (anonymous):

help me plz

OpenStudy (shamil98):

Simplify it to get your answer

OpenStudy (shamil98):

\[\Huge \frac{ 4(5) }{ 2(4) - 6 }\] multiply and reduce

OpenStudy (anonymous):

4x5=20

OpenStudy (anonymous):

wait a min

OpenStudy (anonymous):

4x4=16 2x5=10

OpenStudy (anonymous):

can you help me out

OpenStudy (anonymous):

im confussed

OpenStudy (shamil98):

\[\huge \frac{ 4 \times 5 }{ (2 \times 4) - 6 }\] \[\huge \frac{ 20 }{ 8-6 } = \frac{ 20 }{ 2 }= 10\]

OpenStudy (anonymous):

it 2 right

OpenStudy (anonymous):

no 1o